# The High W Challenge: Robust Neutrino Energy Estimators for LArTPCs

### v1

@@ Abstract
Accurate determination of the neutrino energy is central to precision oscillation measurements. In this work, we introduce the W2-based estimator, a new neutrino energy estimator based on the measurement of the final-state hadronic invatiant mass. This estimator is particularly designed to be employed in liquid-argon time-projection chambers exposed to broadband beams that span the challenging transition region between shallow inelastic scattering and deep inelastic scattering. The performance of the W2-based estimator is compared against four other commonly used estimators. The impact of the estimator choice is evaluated by performing measurements of $\delta_{CP}$ and $\Delta m^{2}_{23}$ in a toy long-baseline oscillation analysis. We find that the W2-based estimator shows the smallest bias as a funciton of true neutrino energy and it is particularly stable against the mismodelling of lepton scattering angle, missing energy, hadronic invariant mass and final state interactions. Such an inclusive channel complements well the strength of more exclusive methods that optimizes the energy resolution. By providing a detailed analysis of strengths, weaknesses and domain of applicability of each estimator, this work informs the combined used of energy estimators in any future LArTPC-based oscillation analysis.
@@ I Introduction
Neutrino oscillations are the only evidence of beyond the Standard Model (BSM) reproducible in the laboratory and provide a unique window into unresolved questions in neutrino physics, such as the ordering of neutrino masses and the potential for CP violation in the lepton sector [T2K:2019bcf].
@@
In a simplified but pedagogically useful scenario, the oscillations probability between two neutrinos of different flavors can be expressed as:
@@
$P(\nu_{\alpha}\rightarrow\nu_{\beta})=\sin^{2}(2\theta)\sin^{2}\left(1.27\frac{\Delta m^{2}_{21}\;L\,[\mathrm{km}]}{E_{\rm true}\,[\mathrm{GeV}]}\right),$ (1)
@@
where $L$ is the baseline (source-to-detector distance), $E_{\rm true}$ is the true neutrino energy, $\theta$ is the mixing angle, and $\Delta m^{2}_{21}$ is the squared mass difference between the two neutrino mass eigenstates. While the oscillation probability is a function of the true neutrino energy, $E_{\rm true}$, detectors can only measure visible final state quantities to estimate the neutrino energy, $E_{\rm est}$. The relationship between these quantities is nontrivial, reflecting the combined effects of the incident neutrino flux and of neutrino–nucleus interactions even before accounting for detector response.
@@
Current and future accelerator-based neutrino experiments such as the Short-Baseline Neutrino (SBN) [Machado:2019oxb] program and the Deep Underground Neutrino Experiment (DUNE) [DUNE:2015lol, DUNE:2016hlj] benefit from the use of broadband energy beams, which enable the sampling of multiple oscillation maxima. This approach enhances sensitivity to oscillation parameters, including CP-violating phases and the neutrino mass ordering. Its drawback, however, is that a wide neutrino energy spectrum inevitably samples a broad range of neutrino–nucleus interaction channels. These span from charged-current quasi-elastic (CCQE) scattering at lower energies, through baryonic resonance production (RES) and meson-exchange current (MEC) processes, up to deep inelastic scattering (DIS) at higher energies. The overlapping nature of these regimes complicates both the modeling of interaction dynamics and the reconstruction of the incident neutrino energy, which is essential for precise oscillation analyses. While previous work has focused on exclusive channels [Furmanski:2016wqo] or on advanced machine learning reconstruction techniques [Kopp:2024lch], we are interested in exploring more inclusive channels, which, by definition, improve of the sample statistics. We assume that ongoing progress in LArTPC reconstruction algorithms will eventually enable high-fidelity reconstruction of most events. Under these conditions, more inclusive analyses may prove to be not only statistically advantageous but also resilient to model and detector-related uncertainties.
@@
This paper explores a new approach to analyzing the final state kinematics, based on the calculation of the final state hadronic invariant mass. This method targets high hadron multiplicity events expected from the interactions of multi-GeV neutrinos in future LArTPCs. This method is then benchmarked against several established methods of estimating neutrino energy, introduced in Section III. Since we lack direct access to nature’s true interaction dynamics, we rely on several event generators, each of which embodies different theoretical assumptions to probe different kinematic scenarios as a function of the true energy. The different event generators employed in this study are introduced in Section II, together with the adopted reconstruction assumptions. We compare the performance of the different estimators by first calculating their bias and variance with respect to neutrino energy in Section V, and then with respect to other kinematic variables in Section VI. Finally, we assess their impact on oscillation measurements through a series of tests presented in Sections VII and VIII.
@@ II Simulation of Neutrino Events
The evaluation of neutrino energy estimators necessarily relies on a set of working assumptions. As a case study, we adopt the neutrino flux provided by the DUNE Collaboration [DUNE:2015lol], without including the effects of oscillation outside of the studies presented in Sections VII and  VIII. We restrict the true neutrino energies to the range 0.5–6 GeV, which is most relevant for oscillation measurements. Any comparable flux that covers the shallow to deep inelastic transition region would serve equally well for the purposes of this study. The neutrino–nucleus interaction models are listed in Section II.1, and our idealized particle reconstruction is outlined in Section II.2. These ingredients provide a controlled environment in which to test and compare the five different estimators of the neutrino energy.
@@ II.1 Event generators
  • GENIE version 3.6.2 [Andreopoulos:2009rq], from which we use the G18_10a_02_11a tune.
  • NuWro 25.03.1 [NuWro:UserGuide].
  • NEUT version 5.8.0 [Hayato:2002sd].
  • GiBUU 2025 [Buss:2011mx].
@@ II.2 Reconstruction Assumptions
In this work, we assume an idealized particle reconstruction accuracy, in which the detection of final-state particles is taken to be 100% efficient above the set of realistic detector thresholds reported in Table 1. The adopted thresholds are consistent with literature [Friedland:2018vry] and motivated by expected capabilities for future LArTPCs [DUNE:2015lol]. In all calculations, we neglect rare final state particles such as kaons and hyperons and we assume neutrons are not detected.
Table 1: Detection thresholds for perfect reconstruction, with values corresponding to the momentum energy of each particle.
@@
Unless otherwise stated, we assume perfect kinematic reconstruction of particles above the thresholds in Table 1. When discussing the impact of detector resolution effects on neutrino energy estimation, we assume the momentum resolutions shown in Table. 2. These are estimated from the performance of the MicroBooNE detector on fully contained particles [MicroBooNE:2025pxk, MicroBooNE:2024bnl, MicroBooNE:2025phj, MicroBooNE:2022tdd]. Additionally, when considering the calorimetric method introduced in Section III.4, we assume a resolution on the total energy of 20%. This choice is justified by the recent MicroBooNE results in Ref. [MicroBooNE:2019efx, MicroBooNE:2022tdd].
Table 2: Momentum resolutions for the measurements of different final state particles used to estimate neutrino energy. When applying the total visible energy method, we also assume an overall energy resolution of 20%.
@@ III Energy Estimators
We compare five energy estimators: (i) one that assumes all interactions are CCQE-like (CCQE-like estimator); (ii) a new approach that improves on the CCQE-like estimator by accounting for multiproton knockout and pion-producing interactions through the hadronic invariant mass ($W^{2}$ estimator); (iii) the sum of the lepton energy and proton kinetic energies (proton-based estimator); (iv) the total energy of all final-state particles, intended to mimic calorimetric reconstruction (calorimetric method); and (v) the Sobczyk–Furmanski method [Furmanski:2016wqo] (SF Estimator).
@@ III.1 CCQE-like Estimator
This estimator assumes charged-current quasi-elastic (CCQE) kinematics to reconstruct neutrino energies, predominantly using the lepton momentum and deflection angle with respect to the beam direction. The estimated neutrino energy ($E^{\rm CCQE}_{\rm est}$) is defined as:
@@
where $m_{p}$, $m_{n}$, and $m_{l}$ are the masses of the proton, neutron, and scattered charged lepton, respectively, $E_{b}$ is a binding energy correction which we take to be 30 MeV for argon, and $E_{l}$, $p_{l}$,and $\theta_{l}$ are the energy, momentum, and scattering angle of the charged lepton.
@@ III.2 W2-based Estimator
The W2-based estimator improves on the CCQE-like method through utilization of the visible hadronic invariant mass($W_{\rm vis}$) to expand the methods assumptions and include meson producing interactions. The initial state mass is adjusted according to the number of visible final state protons to account for the existence of multi-nucleon knockout interactions. The expression for the estimated energy $\left(E^{W^{2}}_{\rm est}\right)$ is:
$E^{W^{2}}_{\rm est}=\frac{W_{\rm vis}^{2}-n_{p}^{2}(m_{n}-E_{b})^{2}-m_{l}^{2}+2n_{p}(m_{n}-E_{b})E_{l}}{2(n_{p}m_{n}-n_{p}E_{b}-E_{l}+p_{l}\cos\theta)},$ (3)
where $n_{p}$ is the number of visible protons in the final state and $W_{\rm vis}$ is the visible hadronic invariant mass. This method is only applied to final states with at least one visible proton.
@@ III.3 Proton-Based Estimator
Designed specifically for low-energy neutrino interactions, this estimator combines detailed track reconstruction with calorimetric measurements. The expression for the estimated energy ($E^{\rm PB}_{\rm est}$) in this instance is given by the lepton total energy ($E_{l}$) plus the sum of the kinetic energies of all visible protons ($E_{k}$):
$E^{\rm PB}_{\rm est}=E_{l}+\sum_{\rm protons}E_{k}.$ (4)
Similar to the method used in the MicroBooNE analyses testing the MiniBooNE anomaly [MicroBooNE:2021wad, MicroBooNECollaboration:2024cpi], this estimator is only be applied to the exclusive topologies containing no visible mesons in the final state.
@@ III.4 Calorimetric Method
Experiments relying on the detection of exclusively DIS events, such as IceCube, typically employ calorimetry or quantities analogous to calorimetric information. This method is distinguished from the others in this review in that high level reconstructed objects such as individual particle kinematics are not required; the energy measurement can be performed in a LArTPC by measuring the total charge collected by the readout.
In this case, the neutrino energy is estimated by summing the measured muon and hadronic energy deposits. Thus, we model the behavior of this estimator by summing the total energy of the lepton, the kinetic energies of all visible protons, and the total energies of all visible pions. The expression for the estimated energy ($E^{\rm SKE}_{\rm est}$) is thus given by the lepton total energy ($E_{l}$) plus the sum of the kinetic energies of all visible protons ($E_{k}$) and pions ($E_{\pi}$):
$E^{\rm SKE}_{\rm est}=E_{l}+\sum_{\rm protons}E_{k}+\sum_{\rm pions}E_{\pi}.$ (5)
Unlike the other methods explored in this paper, this method does not rely on high level reconstructed objects such as particle tracks or electromagnetic showers.
@@ III.5 Sobczyk-Furmanski Estimator
The Sobczyk-Furmanski (SF) method [Furmanski:2016wqo] analyses the energy balance by separating the kinematics into transverse (T) and longitudinal (L) components, to infer the additional energy supplied by the motion of the struck nucleon. The neutrino energy is expressed in terms of the mass of the parent nucleus, $M_{A}$, the mass of the daughter nucleus $M_{A-1}$, and the combined transverse/longitudinal momenta of the lepton-proton system, $p^{T,L}=p^{T,L}_{\rm prot}+p^{T,L}_{l}$. The estimated neutrino energy ($E^{\rm SF}_{\rm est}$) is then obtained from the following expressions:
@@ III.6 Inclusivity Considerations
An important factor in comparing different energy estimation techniques is the degree of inclusiveness of the event sample to which each method can be validly applied, as this directly impacts the available statistics. The CCQE-like and SKE estimators are valid for all hadronic final states, though the CCQE-like estimator is not expected for perform well on events that are not true CCQE. The $W^{2}$ and proton-based estimators both require at least one visible proton, with the latter additionally requiring the absence of visible pions. The SF estimator is the most exclusive, requiring a clean topology with exactly one visible proton and no pions.

### v2

@@ Abstract
Accurate determination of the neutrino energy is central to precision oscillation measurements. In this work, we introduce the W2-based estimator, a new neutrino energy estimator based on the measurement of the final-state hadronic invariant mass. This estimator is particularly designed to be employed in liquid-argon time-projection chambers exposed to broadband beams that span the challenging transition region between shallow inelastic scattering and deep inelastic scattering. The performance of the W2-based estimator is compared to four other commonly used estimators. The impact of the estimator choice is evaluated by performing measurements of $\delta_{CP}$ and $\Delta m^{2}_{23}$ in a toy long-baseline oscillation analysis. We find that the W2-based estimator shows the smallest bias as a function of true neutrino energy and it is particularly stable against the mismodelling of lepton scattering angle and momentum, missing energy, hadronic invariant mass and final state interactions. However, studies of the resolution of each estimator as a function of true neutrino energy show this is somewhat offset by worse energy resolution when perfect modeling of these quantities is assumed. This estimator is valid for events with at least one proton and any number of pions; an inclusive channel that complements the strength of more exclusive methods that optimize the energy resolution. By providing a detailed analysis of the strengths, weaknesses and domain of applicability of each estimator, this work informs the combined use of energy estimators in any future LArTPC-based oscillation analysis.
@@ I Introduction
Neutrino oscillations are the only evidence of physics beyond the Standard Model (BSM) reproducible in the laboratory and provide a unique window into unresolved questions in neutrino physics, such as the ordering of neutrino masses and the potential for CP violation in the lepton sector [2].
@@
In a simplified but pedagogically useful scenario, the oscillation probability between two neutrinos of different flavors can be expressed as:
@@
$P(\nu_{\alpha}\rightarrow\nu_{\beta})=\sin^{2}(2\theta)\sin^{2}\left(1.27\frac{\Delta m^{2}\;L\,[\mathrm{km}]}{E_{\rm true}\,[\mathrm{GeV}]}\right),$ (1)
@@
where $L$ is the baseline (source-to-detector distance), $E_{\rm true}$ is the true neutrino energy, $\theta$ is the mixing angle, and $\Delta m^{2}$ is the squared mass difference between the two neutrino mass eigenstates. While the oscillation probability is a function of the true neutrino energy, $E_{\rm true}$, detectors can only measure visible final state quantities to estimate the neutrino energy, $E_{\rm est}$. The relationship between these quantities is non-trivial and depends on the neutrino–nucleus interaction processes involved even before accounting for detector response.
@@
Current and future accelerator-based neutrino experiments such as the Short-Baseline Neutrino (SBN) [10, 37] program and the Deep Underground Neutrino Experiment (DUNE) [11, 12] are placed on-axis in broadband energy beams, which enable sampling of very wide energy ranges and probing of multiple oscillation maxima. This approach enhances sensitivity to oscillation parameters, including CP-violating phases and the neutrino mass ordering. Its drawback, however, is that a wide neutrino energy spectrum inevitably samples a broad range of neutrino–nucleus interaction channels. These span from charged-current quasi-elastic (CCQE) scattering at lower energies, through baryonic resonance production (RES) and meson-exchange current (MEC) processes, up to deep inelastic scattering (DIS) at higher energies. The overlapping nature of these regimes complicates both the modeling of interaction dynamics and the reconstruction of the incident neutrino energy, which is essential for precise oscillation analyses. While previous work has focused on exclusive channels [28], calorimetry [18, 25] or on advanced machine learning reconstruction techniques [34], we are interested in exploring more inclusive channels, which, by definition, improve the sample statistics. We assume that ongoing progress in LArTPC reconstruction algorithms will eventually enable high-fidelity reconstruction of most events. Under these conditions, more inclusive analyses may prove to be not only statistically advantageous, but also more resilient to model and detector-related uncertainties.
@@
This paper explores a new approach to analyzing the final state kinematics, based on the calculation of the final state hadronic invariant mass. This method targets high hadron multiplicity events expected from the interactions of multi-GeV neutrinos in future LArTPCs. This method is then benchmarked against several established methods of estimating neutrino energy, introduced in Section III.
To model the complex neutrino-nucleus interactions, we rely on several event generators, each of which embodies different theoretical assumptions to probe different kinematic scenarios as a function of the true energy. The event generators employed in this study are introduced in Section II, together with the adopted reconstruction assumptions. We compare the performance of the different estimators by first calculating their bias and variance with respect to neutrino energy in Section V, and then with respect to other kinematic variables in Section VI. Finally, we assess their impact on oscillation measurements through a series of tests presented in Sections VII and VIII.
@@ II Simulation of Neutrino Events
The evaluation of neutrino energy estimators necessarily relies on a set of working assumptions. As a case study, we adopt the neutrino flux provided by the DUNE Collaboration [11]. We do not include the effects of oscillation outside of the studies presented in Sections VII and  VIII. We restrict the true neutrino energies to the range 0.5–6 GeV, which is most relevant for oscillation measurements, and considerably more energetic than the beams used in existing LArTPC experiments [10]. Any comparable flux that covers the shallow to deep inelastic transition region would serve equally well for the purposes of this study. The neutrino interaction event generators are listed in Section II.1, and our idealized particle reconstruction is outlined in Section II.2. These ingredients provide a controlled environment in which to test and compare the five different neutrino energy estimators.
@@ II.1 Event generators
  • GENIE version 3.6.2 [17], from which we use the G18_10a_02_11a tune. The tune employs the Nieves model of quasielastic interactions [39] with the BBBA05 vector form factor parameterization [22] and including random phase approximation (RPA) corrections. The Beger-Sehgal for resonance production [41, 20] and Valencia model of CC meson exchange current [45, 40, 30]. Resonances up to $W=1.9$ GeV are simulated after which the generator transitions to the Bodek-Yang DIS model [21]. Hadronization is performed with the AGKY model [48]. Final state interactions are simulated using the hA2018 model.
  • NuWro 25.03.1 [46]. NuWro employs the LLwelyn-Smith model [36] of quasielastic interactions augmented with RPA corrections, BBBA05 vector form factors [22] and the MINERvA parameterisation [24] of the axial form factor. Meson exchange current is simulated with Valencia model [45, 40, 30], and the Ghent hybrid model of single pion production [47]. Unlike GENIE, NuWro does not simulate higher resonances and instead switches to modeling the interaction as DIS when the invariant mass of the W boson-nucleon system exceeds 1.5 GeV. DIS interactions are simulated with the Bodek-Yang model [21] with hadronization routines outsourced to Pythia 6 [43]. Final state interactions are modeled with a cascade simulation developed in-house, the most complete description of which is found in [29].
  • NEUT version 5.8.0 [31, 32]. This generator uses the Llwelyn-Smith model [36] for CCQE interactions. Meson exchange current is simulated with the Valencia model [45, 40, 30], and single pion production using the Berger-Sehgal model [20]. Mutli-pion production events are modeled using a series of fits to hadron multiplicity data from bubble chambers [26, 33] before the model switches to deep inelastic scattering at $W=2.0$ GeV, simulated with Pythia 5 [44].
  • GiBUU 2025 [23]. This is distinct from the other three generators in that the neutrino-nucleus scattering cross sections are calculated through solution of nuclear transport equations, with single nucleon cross section calculations detailed in  [23], excluding DIS which is modeled using Pythia [43]
@@ II.2 Reconstruction Assumptions
In this work, we assume an idealized particle identification and reconstruction accuracy, in which the detection of final-state particles is taken to be 100% efficient above the set of realistic detector thresholds reported in Table 1, and there is no misidentification of one particle type as another. The adopted thresholds are consistent with literature [27] and motivated by expected capabilities for future LArTPCs [11]. In all calculations, we assume neutrons are not detected. We also ignore flavored hadrons such as kaons and hyperons as their impact on the estimated energy spectra will be negligible on account of their rarity and the detection performance of LArTPCs on these particles is not yet well understood.
Table 1: Detection thresholds for perfect reconstruction, with values corresponding to the momentum of each particle.
@@
Unless otherwise stated, we assume perfect kinematic reconstruction of particles above the thresholds in Table 1. When discussing the impact of detector resolution effects on neutrino energy estimation, we assume the momentum resolutions shown in Table. 2. These are estimated from the performance of the MicroBooNE detector on fully contained particles [8, 6, 7, 3]. Additionally, when considering the calorimetric method introduced in Section III.4, we assume a resolution on the total energy of 20%. This choice is justified by the recent MicroBooNE results in Ref. [15, 3].
Table 2: Momentum resolutions for the measurements of different final state particles used to estimate neutrino energy. When applying the calorimetric method, we also include a separate calculation using an overall energy resolution of 20%.
@@ III Energy Estimators
We compare five energy estimators: (i) one that assumes all interactions are CCQE-like (CCQE-like estimator); (ii) a new approach that improves on the CCQE-like estimator by accounting for multiproton knockout and pion-producing interactions through the hadronic invariant mass ($W^{2}$ estimator); (iii) the sum of the lepton energy and proton kinetic energies (proton-based estimator); (iv) the total energy of all final-state particles, intended to mimic calorimetric reconstruction (calorimetric method); and (v) the Sobczyk–Furmanski method [28] (SF Estimator).
@@ III.1 CCQE-like Estimator
This estimator assumes charged-current quasi-elastic (CCQE) kinematics to reconstruct neutrino energies, predominantly using the lepton momentum and deflection angle with respect to the beam direction. The estimated neutrino energy ($E^{\rm CCQE}_{\rm est}$) is defined as [35]:
@@
[…]
@@ III.2 W2-based Estimator
The W2-based estimator improves on the CCQE-like method through utilization of the visible hadronic invariant mass:
$W^{2}_{\rm vis}=\left(\sum_{i}E_{i}\right)^{2}-\left(\sum_{i}\textbf{p}_{i}\right)^{2},$ (3)
where $E_{i}$ and $\textbf{p}_{i}$ are the total energy and 3-momentum of the $i$th visible final state particle respectively, and initial state mass is adjusted according to the number of visible final state protons ($n_{p}$). These modifications extend assumptions made by the CCQE-like method to account for the existence of meson producing and multi-nucleon knockout interactions. The expression for the estimated energy $\left(E^{W^{2}}_{\rm est}\right)$ is:
$E^{W^{2}}_{\rm est}=\frac{W_{\rm vis}^{2}-n_{p}^{2}(m_{n}-E_{b})^{2}-m_{l}^{2}+2n_{p}(m_{n}-E_{b})E_{l}}{2(n_{p}m_{n}-n_{p}E_{b}-E_{l}+p_{l}\cos\theta_{l})}.$ (4)
This method is only applied to final states with at least one visible proton.
@@ III.3 Proton-Based Estimator
Designed specifically for low-energy neutrino interactions, this estimator combines detailed track reconstruction with calorimetric measurements. The expression for the estimated energy ($E^{\rm PB}_{\rm est}$) in this instance is given by the lepton total energy ($E_{l}$) plus the sum of the kinetic energies of all visible protons ($T_{p}$):
$E^{\rm PB}_{\rm est}=E_{l}+\sum_{\rm protons}T_{p}.$ (5)
Similar to the method used in the MicroBooNE analyses testing the MiniBooNE anomaly [5, 9], this estimator is only be applied to the exclusive topologies containing no visible mesons in the final state.
@@ III.4 Calorimetric Method
Some other tracking calorimeter detectors studying multi-GeV neutrinos, such as NOvA [19] and MINERvA [16], rely methods driven by calorimetry to measure the energy of the final state hadronic system [13, 42], without attempting to identify individual particles. Furthermore, detectors focused on even very high energies such as IceCube [14] detect exclusively DIS interactions, and are purely calorimetric (or measure quantities analogous to energy deposition). This method is distinguished from the others in this review in that high level reconstructed objects such as individual particle kinematics are not required; the energy measurement can be performed in a LArTPC by measuring the total charge collected by the readout that is not attributed to cosmic rays or radiological backgrounds.
In this case the estimated energy ($E^{\rm SKE}_{\rm est}$) is modeled by summing the total lepton total energy ($E_{l}$), the kinetic energies of all visible protons ($T_{p}$), and total energies of all pions ($E_{\pi}$):
$E^{\rm SKE}_{\rm est}=E_{l}+\sum_{\rm protons}T_{p}+\sum_{\rm pions}E_{\pi}.$ (6)
Unlike the other methods explored in this paper, this method does not rely on high level reconstructed objects such as particle tracks or showers.
@@ III.5 Sobczyk-Furmanski Estimator
The Sobczyk-Furmanski (SF) method [28] analyses the energy balance by separating the kinematics into transverse (T) and longitudinal (L) components, to infer the additional energy supplied by the motion of the struck nucleon. The neutrino energy is expressed in terms of the mass of the parent nucleus, $M_{A}$, the mass of the daughter nucleus $M_{A-1}$, and the combined transverse/longitudinal momenta of the lepton-proton system, $p^{T,L}=p^{T,L}_{\rm prot}+p^{T,L}_{l}$. The estimated neutrino energy ($E^{\rm SF}_{\rm est}$) is then obtained from the following expressions:
@@ III.6 Inclusivity Considerations
An important factor in comparing different energy estimation techniques is the degree of inclusiveness of the event sample to which each method can be validly applied, as this directly impacts the available statistics. The CCQE-like and calorimetric estimators are valid for all hadronic final states, though the CCQE-like estimator is not expected for perform well on events that are not true CCQE, hence the two modified forms. The $W^{2}$ estimator both requires at least one visible proton, and the proton based estimator events with no visible pions. The SF estimator is the most exclusive, requiring a clean topology with exactly one visible proton and no pions.
@@
(a) Rates normalized to equal arbitrary exposure.
(b) Fractions of samples used by each exclusive channel.